Search for Single Top Production

Pekka Sinervo University of Toronto

Overview

• Standard Model Framework for Top • Single Top Production

• Search in Run I Data – Technique

– Results

• Prognostications for Run II

Work done primarily by Sarah Wolinski and Pierre Savard

Top Quark Production

Top has been observed through the LO SM process

pp̄ → tt̄

σ_{t t} = 6.5_{−1.4}^{+1.7} pb

• Theory predictions range from 4.7 to 5.5 pb However, this is not the only source of top quarks!

Single Top Production

Single top production is a much lower rate process

• Sensitive to Γ_{t} and V_{tb}

\frac{|V_{tb}|^{2}}{|V_{td}|^{2} + |V_{ts}|^{2} + |V_{tb}|^{2}} = 0.99 ± 0.29

• Also could have contributions from non-SM sources

There are two primary SM production processes

1. t-channel W − g fusion ⇒ W^{+} bq final state

2. s-channel W^{∗} process ⇒ W^{+} bb̄ final state

• Predicted cross sections are 1.7 ± 0.3 and 0.7 ± 0.1 pb CAP Congress, 7 Jun 2000

Search Technique

Kinematics of production processes different

• Use the same selection for both processes – 1, 2 or 3 jets with E_{T} > 15 Gev and |η| < 2.0

– At least one b-tagged jet

– Require 140 ≤ M_{top} ≤ 210 GeV/c^{2}

• Fit H_{T} distribution with signal and background terms

H_{T} ≡ ∑_{partons} E_{T}

W^{∗}  W -gluon
σ_{theory} (pb)  0.73 ± 0.10  1.7 ± 0.3
Integrated luminosity (pb^{−1})  106 ± 4  106 ± 4
W + n jet acceptance (%)  4.5 ± 0.7  6.7 ± 1.0
SVX b-tag efficiency (%)  37.5 ± 3.1  24.8 ± 2.0
Number of expected events  1.3 ± 0.3  3.0 ± 0.6

Top Mass Reconstruction

CAP Congress, 7 Jun 2000

H_{T} Reconstruction

The variable H_{T} is an effective discriminator

• QCD Background

• t t background

• Single top signal

Results of Search

Observed 65 events in this sample

• Largest single background are mistags

– tt̄ itself is a problem now

Process  Event Yield
Mistags  25.9 ± 7.5
W bb̄ + W cc̄  28.7 ± 7.5
W c  17.2 ± 6.2
Non-W  12.1 ± 4.1
Miscellaneous  5.1 ± 0.9
After M_{t} cut  54 ± 11
tt̄  8.4 ± 2.7
W -gluon  3.0 ± 0.6
W^{∗}  1.3 ± 0.3
Total expected  67 ± 11
Observed  65

Not yet sensitive to SM single top rate

• Can set limits on total single top rate

Single Top Production Cross Section Limits

Use H_{T} distribution to fit signal and background shapes

• Set upper limit using a Bayesian technique

• Fit gives σ_{t} < 13.5 pb at 95% CL

Still well away from SM rate

• Can conclude that there is no huge anomaly

• It is a feasable measurement for Run II

– Can get S/N = 2.9 on W + 2, 3 jet sample CAP Congress, 7 Jun 2000

Run II Forecasts

Run II will start this summer at √s = 2.0 TeV

• CDF will begin with engineering run (no SVX II)

• Both detectors will be completed by end of 2000 • Physics running will commence in 2001

Will be worth the wait

• Begin with ∼ 3 fb^{−1}/year per detector

• Finish up with ∼ 5 fb^{−1}/year

• Literal “Top Factory” with – 300 dilepton events

– 3000 tagged lepton+jet events – 1500 events in “mass sample”

Single top production

• Measure σ_{t} with 10% uncertainty

• Measure V_{tb} with 12% uncertainty

Summary

Debut of top using Run I data coming to close

• Have seen that it looks much like a SM fermion • Clearly observed through QCD tt̄ production

• Single top harder to extract – 95% CL upper limit of σ_{t} < 13.5 pb at 95% CL

All this sets the stage for Run II

• Run for 2-4 years, with goal of ∼ 6 fb^{−1} ⇒ Increase top quark sample ×50

– Should be able to measure single top production ∗ S/N ∼ 2

∗ σ_{t} ∼ 10%